# SAGMAN: Stability Analysis of Graph Neural Networks on the Manifolds

### v3

@@ 3.5 Phase 3: Stability Analysis on the Manifolds
DMD with Resistance Distance Metric.
Let $M$ denote the mapping function of an ML model that operates on input $X$ to yield output $Y$, i.e., $Y=M(X)$. The effective-resistance distances between data samples $p$ and $q$ on the input and output graph-based manifolds can be computed by $d^{eff}_{X}(p,q)=e_{p,q}^{\top}L_{X}^{+}e_{p,q}$, and $d^{eff}_{Y}(p,q)=e_{p,q}^{\top}L_{Y}^{+}e_{p,q}$, respectively, where $L_{X}^{+}$ and $L_{Y}^{+}$ denote the Moore–Penrose pseudoinverses of graph Laplacian matrices of the input and output graph-based manifolds, respectively. For any arbitrary node pair $p$ and $q$ on graph-based manifolds, the maximum DMD measured using resistance-distance metric can be obtained by solving the following combinatorial optimization problem (Cheng et al., 2021):
$\delta_{max}^{M}(p,q)=\mathop{\max_{\forall p,q\in V}}_{p\neq q}\frac{d^{eff}_{Y}(p,q)}{d^{eff}_{X}(p,q)}=\mathop{\max_{\forall p,q\in V}}_{p\neq q}\frac{e_{p,q}^{\top}L_{Y}^{+}e_{p,q}}{e_{p,q}^{\top}L_{X}^{+}e_{p,q}}.$ (9)
Then the best Lipschitz constant $K^{*}$ of the mapping function $M$ is upper bounded by the largest eigenvalue of $L_{Y}^{+}L_{X}$ according to Lemma 2.2 (Cheng et al., 2021):
$\lambda_{max}(L_{Y}^{+}L_{X})\geq K^{*}\geq{\delta_{max}^{M}(p,q)}.$ (10)
@@
Recent theoretical advancements (Cheng et al., 2021) motivate us to exploit the largest generalized eigenvalues and their corresponding eigenvectors for the stability assessment of individual nodes in GNNs. Specifically, we first compute the weighted eigensubspace matrix $V_{s}\in{\mathbb{R}^{|V|\times s}}$ for spectral embedding on the input graph-based manifold $G_{X}=(V,E_{X})$ with $V$ nodes: $V_{s}\overset{\mathrm{def}}{=}\left[{v_{1}}{\sqrt{\zeta_{1}}},...,{v_{s}}{\sqrt{\zeta_{s}}}\right],$ where $\zeta_{1},\zeta_{2},...,\zeta_{s}$ represent the first $s$ largest eigenvalues of $L_{Y}^{+}L_{X}$ and $v_{1},v_{2},...,v_{s}$ are the corresponding eigenvectors. Subsequently, $G_{X}$ is embedded using $V_{s}$, where each node is represented by an $s$-dimensional embedding vector. The stability of an edge $(p,q)\in E_{X}$ can be then estimated by computing the spectral embedding distance between two end nodes $p$ and $q$. The stability score of a node $p$ can be further computed as follows:
$\begin{split}\textbf{score}(p)\overset{\mathrm{def}}{=}\frac{1}{|\mathbb{N}_{X}(p)|}\sum_{q_{i}\in\mathbb{N}_{X}(p)}\left(\|V_{s}^{\top}e_{p,q_{i}}\|_{2}^{2}\right)\\ \propto\frac{1}{|\mathbb{N}_{X}(p)|}\sum_{q_{i}\in\mathbb{N}_{X}(p)}\left(\delta^{M}(p,q_{i})\right)^{3}\end{split}$ (11)
@@
where $\mathbb{N}_{X}(p)$ represents the set of neighbors of node $p$ in the input graph-based manifold $G_{X}$. This node stability score can effectively serve as a surrogate for the local Lipschitz constant, which is analogous to $\|\nabla_{X}F(p)\|$ under the manifold setting (Cheng et al., 2021). For a more detailed exposition, please refer to Appendix F.
3.6 Complexity of SAGMAN
@@
The detailed SAGMAN flow is shown in Appendix B. For spectral graph embedding, we can exploit fast multilevel eigensolvers that allow computing the first $c$ Laplacian eigenvectors in nearly-linear time $O(c|V|)$ without loss of accuracy (Zhao et al., 2021). The $k$-nearest neighbor algorithm (Malkov & Yashunin, 2018) has a nearly-linear computational complexity $O(|V|\log|V|)$. Spectral sparsification via LRD decomposition has $O(|V|dm)$ time complexity. By leveraging fast generalized eigensolvers (Koutis et al., 2010; Cucuringu et al., 2016) all DMD values can be computed within $O(|E|)$ time, where the $V$/$E$ denotes the number of nodes/edges, $d$ denotes the average degree of the matrix, and $m$ is the order of Krylov subspace. The SAGMAN is capable of handling very large datasets due to its nearly-linear time complexity. For instance, SAGMAN takes $391.15$ seconds to process the “ogbn-arxiv” dataset with $169,343$ nodes and $1,166,243$ edges on a system with an $i9$ $10900kf$ processor, $32GB$ RAM, and an NVIDIA Titan RTX $24GB$ GPU.
@@ 4 Experiments
In this paper, We conduct a comparison between exact resistance distances from the original graph and their approximations derived from the constructed manifold. Then, we present a series of fundamental numerical experiments designed to showcase the effectiveness of our proposed metric in quantifying the stability of GNNs under various perturbations in node features and edges. Additionally, we extend our analysis to the realm of recommendation systems, demonstrating how our metric measures their stability effectively. Furthermore, we highlight the efficacy and efficiency of our SAGMAN-guided approach in executing graph adversarial attacks. Then, we show how to significantly enhance the stability of GNN models leveraging the low-dimensional graph-based input manifold created by SAGMAN. Lastly, We present the runtime performance of SAGMAN across various graph scales.
Experimental Setup. See Appendix E for detailed descriptions of all graph datasets used in this work. We employ the most popular backbone GNN models including GCN (Kipf & Welling, 2016), GPRGNN (Chien et al., 2020), GAT (Veličković et al., 2017), APPNP (Gasteiger et al., 2018), and ChebNet (Defferrard et al., 2016). The recommendation system is based on PinSage (Ying et al., 2018). Perturbations include Gaussian noise evasion attacks and adversarial attacks (DICE (Waniek et al., 2018), Nettack (Zügner et al., 2018), and FGA (Chen et al., 2018)). The input graph-based manifolds are constructed using graph adjacency and node features, while the output graph-based manifold is created using post-softmax vectors. To showcase SAGMAN’s effectiveness in differentiating stable from unstable nodes, we apply SAGMAN to select $1\%$ of the entire dataset as stable nodes, and another $1\%$ as unstable nodes. This decision stems from that only a portion of the dataset significantly impacts model stability (Cheng et al., 2021; Hua et al., 2021; Chang et al., 2017). Additional evaluation results on the entire dataset can be found in Appendix G. We quantify output perturbations using cosine similarity and Kullback-Leibler divergence (KLD). Additional insights on cosine similarity, KLD, and accuracy can be found in Appendix A. For a large-scale dataset “ogbn-arxiv”, due to its higher output dimensionality, we focus exclusively on accuracy comparisons. This decision is informed by the KLD estimator’s $n^{-\frac{1}{d}}$ convergence rate (Roldán & Parrondo, 2012), where $n$ is the number of samples and $d$ is the dimension. In this paper, our spectral embedding method consistently utilizes the smallest $50$ eigenpairs for all experiments, unless explicitly stated otherwise.
@@ 4.1 Evaluation of Graph Dimension Reduction (GDR)
There is a valid concern about how much the constructed input graph-based manifold might change the original graph’s structure. To this end, we compare the exact resistance distances calculated using the complete set of eigenpairs with approximate ones estimated using the embedding matrix $U_{k}$ in Definition 3.1 where various $k$ values have been considered in Table 1. This empirical evidence shows that using a small number of eigenpairs can effectively approximate the original effective-resistance distances. Table 2 empirically shows the SAGMAN-guided stability analysis with GDR can always distinguish stable and unstable nodes, while the one without GDR can not. Additional related results for various datasets and GNNs have been reported in Appendix C.
@@
Table 2: Nettack adversarial attack targeting SAGMAN-selected Cora samples in GCN. Better results are highlighted in bold.
Nettack Level Cosine Similarities: Stable/Unstable
w/o GDR w/ GDR
1 0.90/0.96 0.99/0.90
2 0.84/0.93 0.98/0.84
3 0.81/0.91 0.97/0.81
@@
Dataset
Nettack’s Recommendation
Nettack/FGA Error Rate
Confidence Ranking
Nettack/FGA Error Rate
SAGMAN
Nettack/FGA Error Rate
@@
Table 4: Error rates for the unstable dataset selected by SAGMAN. ‘Clean’ denotes the original graph without perturbations, while ‘Perturbed’ corresponds to the graph after a Nettack evasion attack. ‘Size’ indicates the fraction of the SAGMAN-selected nodes.
Cora Original Graph Enhanced Graph
Size Clean/Perturbed Clean/Perturbed
1% 0.000/0.791 0.000/0.000
5% 0.008/0.862 0.000/0.000
10% 0.008/0.903 0.008/0.008
15% 0.021/0.927 0.064/0.067
20% 0.024/0.921 0.078/0.086
@@ 4.2 Metrics for GNN Stability Evaluation
We demonstrate the capability of SAGMAN for empirically distinguishing between stable samples and unstable samples in Figure 3. More results regarding different datasets, GNNs, and Nettack attacks are available in Appendix E. It illustrates our metric proves its effectiveness. Last, we report accuracy differences regarding large-scale datasets in Figure 4
Figure 4: SAGMAN-selected stable/unstable samples for the ”ogbn-arxiv” dataset. We report the GCN prediction accuracies under different levels of Gaussian Noise Perturbation $FM+X\eta$, where $FM$ denotes the feature matrix, $\eta$ represents Gaussian noise, and $X$ is the noise perturbation level.
@@ 4.3 Stability of GNN-based Recommendation Systems
Our evaluation employs the PinSage GNN framework (Ying et al., 2018), utilizing the MovieLens 1M dataset (Harper & Konstan, 2015). To construct the input graph, we homogenize various node and edge types into a unified format, followed by calculating the weighted spectral embedding matrix $U_{k}$, which is defined in Definition 3.1. We then extract user-type samples from this matrix to build the low-dimensional input (user) graph-based manifold. To create the output graph-based manifold, we use PinSage’s final output including the top-$10$ recommended items for each user. An output (user) graph-based manifold is subsequently constructed based on Jaccard similarity measures. Table 5 clearly showcases the effectiveness of SAGMAN in differentiating between stable and unstable users.
Table 5: Comparison of the mean Jaccard similarity (MJS) between SAGMAN-selected stable and unstable users at a perturbation level ($l$) ranging from $1$ to $5$ where each selected user is connected to $l$ randomly chosen new items in the input graph. The MJS is computed over $20$ iterations for each perturbation level.
Per. Lev. ($l$) Stable Users Unstable Users
1 0.8513 0.7885
2 0.8837 0.7955
3 0.8295 0.8167
4 0.8480 0.8278
5 0.8473 0.7794
@@ 4.4 SAGMAN-guided Adversarial Targeted Attack
Our GNN training methodology is adapted from the approach delineated by (Jin et al., 2021). We adopt GCN as our base model for Citeseer, Cora, Cora-ml, and Pubmed, with Nettack and FGA serving as the benchmark attack methods. For comparison, target node selection follows Nettack’s recommendation (Zügner et al., 2018), SAGMAN-guided strategies, and a heuristic confidence ranking (Chang et al., 2017). Table 3 shows the error rates after Nettack and FGA attacks. As observed, the SAGMAN-guided attack outperforms both Nettack’s recommendation and confidence ranking, leading to more effective Nettack and FGA attacks. Furthermore, the SAGMAN-guided FGA attack (Chen et al., 2018) also demonstrates superior effectiveness.

### v4

@@ 3.4 Phase 3: Stability Analysis on the Manifolds
In this phase, we perform stability analysis of GNNs by quantifying the distortions between the input and output graph-based manifolds using the Distance Mapping Distortion (DMD) metric.
DMD with Effective Resistance Distance Metric.
Let $M$ be the mapping function of a machine learning model that transforms input data $X$ into output data $Y$, i.e., $Y=M(X)$. To assess the stability of $M$, we employ the DMD metric, which measures how distances between data samples are distorted by $M$. On the input and output graph-based manifolds, we use the effective resistance distance as the metric between nodes. The effective resistance distance between nodes $p$ and $q$ on the input manifold $G_{X}$ is computed as $d^{\mathrm{eff}}_{X}(p,q)=e_{p,q}^{\top}L_{X}^{+}e_{p,q},$ and similarly on the output manifold $G_{Y}$ as $d^{\mathrm{eff}}_{Y}(p,q)=e_{p,q}^{\top}L_{Y}^{+}e_{p,q},$ where $L_{X}^{+}$ and $L_{Y}^{+}$ denote the Moore–Penrose pseudoinverses of the Laplacian matrices $L_{X}$ and $L_{Y}$ of the input and output manifolds, respectively, and $e_{p,q}=e_{p}-e_{q}$ with $e_{p}$ being the standard basis vector corresponding to node $p$. The DMD between nodes $p$ and $q$ is defined as the ratio of the output distance to the input distance:
$\delta^{M}(p,q)=\frac{d^{\mathrm{eff}}_{Y}(p,q)}{d^{\mathrm{eff}}_{X}(p,q)}=\frac{e_{p,q}^{\top}L_{Y}^{+}e_{p,q}}{e_{p,q}^{\top}L_{X}^{+}e_{p,q}}.$ (3)
To quantify the maximum distortion introduced by $M$, we consider the maximum DMD over all pairs of distinct nodes: $\delta_{\mathrm{max}}^{M}$. According to Lemma A.2 in (Cheng et al., 2021), the optimal Lipschitz constant $K^{*}$ of the mapping $M$ is bounded by the largest eigenvalue of $L_{Y}^{+}L_{X}$ and the maximum DMD:
$\delta_{\mathrm{max}}^{M}\leq K^{*}\leq\lambda_{\max}(L_{Y}^{+}L_{X}).$ (4)
This relationship allows us to assess the stability of $M$ using spectral properties of the Laplacian matrices.
@@
Building on the theoretical insights from (Cheng et al., 2021), we utilize the largest generalized eigenvalues and their corresponding eigenvectors of $L_{Y}^{+}L_{X}$ to evaluate the stability of individual nodes in GNNs. We compute the weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ for spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ is the number of nodes. The matrix $V_{s}$ is defined as: $V_{s}=\left[v_{1}\sqrt{\zeta_{1}},\,v_{2}\sqrt{\zeta_{2}},\,\dots,\,v_{s}\sqrt{\zeta_{s}}\right]$, where $\zeta_{1}\geq\zeta_{2}\geq\dots\geq\zeta_{s}$ are the largest $s$ eigenvalues of $L_{Y}^{+}L_{X}$, and $v_{1},v_{2},\dots,v_{s}$ are the corresponding eigenvectors. Using $V_{s}$, we embed the nodes of $G_{X}$ into an $s$-dimensional space by representing each node $p$ with the $p$-th row of $V_{s}$. The stability of an edge $(p,q)\in E_{X}$ can then be estimated by computing the spectral embedding distance between nodes $p$ and $q$: $|V_{s}^{\top}e_{p,q}\|_{2}^{2}$. To assess the stability of individual nodes, we define the stability score of node $p$ as the average embedding distance to its neighbors in the input manifold:
$\text{score}(p)=\frac{1}{|\mathbb{N}_{X}(p)|}\sum_{q\in\mathbb{N}_{X}(p)}\|V_{s}^{\top}e_{p,q}\|_{2}^{2},$ (5)
@@
where $\mathbb{N}_{X}(p)$ denotes the set of neighbors of node $p$ in $G_{X}$. Since $\|V_{s}^{\top}e_{p,q}\|_{2}^{2}$ is proportional to $\left(\delta^{M}(p,q)\right)^{3}$, the node stability score effectively serves as a surrogate for the local Lipschitz constant, analogous to $\|\nabla_{X}M(p)\|$ under the manifold setting (Cheng et al., 2021). For a more detailed derivation and theoretical justification, please refer to Appendix A.11.
3.5 Algorithm Flow of SAGMAN
The Algorithm 1 shows the key steps in SAGMAN.
Algorithm 1 SAGMAN Algorithm
Input: Graph $G=(V,E)$, node features $X$, GNN model
Output: Stability scores for all nodes
  1. 1.
    Compute Spectral Embeddings (GDR):
    • Compute the spectral embedding matrix $U_{k}$ of $G$ using its weighted Laplacian.
  2. 2.
    Augment Node Features:
    • Concatenate $X$ and $U_{k}$ to form the feature matrix $FM=[U_{k},X]$.
  3. 3.
    Construct Input Graph-based Manifold $G_{X}$ (PGM):
    • Build a k-NN graph $G_{\text{dense}}^{X}$ using $FM$.
    • Apply spectral sparsification to $G_{\text{dense}}^{X}$ to obtain $G_{X}$.
  4. 4.
    Apply GNN Model:
    • Compute output representations $Y=\text{GNN}(G_{X},X)$.
  5. 5.
    Construct Output Graph-based Manifold $G_{Y}$ (PGM):
    • Build a k-NN graph $G_{\text{dense}}^{Y}$ using $Y$.
    • Apply spectral sparsification to $G_{\text{dense}}^{Y}$ to obtain $G_{Y}$.
  6. 6.
    Compute Stability Scores (DMD):
    • Compute Laplacians $L_{X}$ and $L_{Y}$ of $G_{X}$ and $G_{Y}$.
    • Solve the generalized eigenvalue problem $L_{Y}V_{k}=\lambda L_{X}V_{k}$ to obtain $V_{k}$.
    • For each node $p\in V$:
      • Compute the stability score: $\text{score}(p)=\frac{1}{|\mathcal{N}_{X}(p)|}\sum_{q\in\mathcal{N}_{X}(p)}\|V_{k}^{\top}(e_{p}-e_{q})\|_{2}^{2}$ where $\mathcal{N}_{X}(p)$ are the neighbors of $p$ in $G_{X}$.
3.6 Time Complexity of SAGMAN
@@
The proposed SAGMAN framework is designed to be efficient and scalable for large graphs. Below, we analyze the time complexity of its key components. We utilize fast multilevel eigensolvers to compute the first $c$ Laplacian eigenvectors. These eigensolvers operate in nearly linear time, $O(c|V|)$, without loss of accuracy (Zhao et al., 2021), where $|V|$ denotes the number of nodes in the graph. To construct the initial graph for manifold learning, we employ the $k$-nearest neighbor algorithm, which has a nearly linear computational complexity of $O(|V|\log|V|)$ (Malkov & Yashunin, 2018). The spectral sparsification step, performed via Low-Resistance-Diameter (LRD) decomposition, has a time complexity of $O(|V|dm)$, where $d$ is the average degree of the graph, and $m$ is the order of the Krylov subspace used in the computation. By leveraging fast generalized eigensolvers (Koutis et al., 2010; Cucuringu et al., 2016), we can compute all DMD values in $O(|E|)$ time, where $|E|$ denotes the number of edges in the graph.
Overall, the dominant terms in the time complexity are nearly linear with respect to the size of the graph. This near-linear scalability allows SAGMAN to handle large graph datasets efficiently. Results for runtime scalability are available in Appendix A.13
@@ 4 Experiments
We validate our proposed SAGMAN framework through comprehensive experiments. First, we compare exact resistance distances from the original graph with their approximations derived from the constructed manifold to demonstrate the manifold’s fidelity. Next, we present numerical experiments showcasing the effectiveness of our metric in quantifying GNN stability. We then highlight the efficacy of our SAGMAN-guided approach in executing graph adversarial attacks. Finally, we demonstrate how leveraging the low-dimensional input manifold created by SAGMAN significantly enhances GNN stability. Details of our experimental setup are provided in Appendix A.5.
@@ 4.1 Evaluation of Graph Dimension Reduction (GDR)
A natural concern arises regarding how much the constructed input graph-based manifold might change the original graph’s structure. To this end, we compare the exact resistance distances calculated using the complete set of eigenpairs with approximate resistance distances estimated using the embedding matrix $U_{k}$ (as defined in Definition 3.1), considering various values of $k$ as shown in Table 1. Our empirical results demonstrate that using a small number of eigenpairs can effectively approximate the original effective-resistance distances. Furthermore, Table 2 shows that the SAGMAN-guided stability analysis with GDR consistently distinguishes between stable and unstable nodes, whereas the analysis without GDR fails to do so. Additional related results for various datasets and GNN architectures are provided in Appendix A.8.
@@
Table 2: Cosine similarities between original and perturbed node embeddings for stable/unstable nodes under Nettack adversarial attacks on the Cora dataset using GCN. Higher cosine similarities for stable nodes and lower for unstable nodes indicate better distinction. Better results are in bold.
Nettack Level 1 2 3
w/o GDR 0.90/0.96 0.84/0.93 0.81/0.91
w/ GDR 0.99/0.90 0.98/0.84 0.97/0.81
@@
Selection Nettack’s default Confidence Ranking SAGMAN
@@
Table 4: Error rates for the Cora and Citeseer datasets before/after Nettack and IG-attack evasion attacks, with and without SAGMAN enhancement.‘Size’ indicates the fraction of the nodes selected. Better results are highlighted in bold.
Nettack Attack Budget Size 1% 5% 10% 15% 20%
Cora w/o SAGMAN 0.00/0.79 0.01/0.86 0.01/0.90 0.02/0.93 0.02/0.92
w/ SAGMAN 0.00/0.00 0.00/0.00 0.01/0.01 0.06/0.06 0.08/0.08
Citeseer w/o SAGMAN 0.00/0.76 0.03/0.82 0.04/0.85 0.04/0.86 0.03/0.85
w/ SAGMAN 0.09/0.42 0.10/0.37 0.12/0.37 0.11/0.35 0.08/0.28
IG-attack Budget Size 1% 5% 10% 15% 20%
Citeseer w/o SAGMAN 0.00/0.76 0.02/0.74 0.02/0.74 0.02/0.72 0.02/0.75
w/ SAGMAN 0.00/0.05 0.02/0.07 0.02/0.07 0.02/0.04 0.02/0.09
Cora w/o SAGMAN 0.00/0.96 0.02/0.83 0.02/0.81 0.02/0.82 0.02/0.81
w/ SAGMAN 0.00/0.00 0.02/0.02 0.02/0.03 0.02/0.02 0.02/0.04
Table 5: Baseline and Perturbed error rates for the Citeseer dataset using GOOD-AT, SAGMAN, and both. Our experiments’ attack method and backbone architecture follow those from Li et al. (2024). The best results are marked in bold, and the second-best results are underlined.
Edge Attack Budget 25 109 219 302 410 550
GCN(baseline) 0.2968 0.3211 0.3590 0.3780 0.4064 0.4301
GOOD-AT 0.2808 0.2820 0.2850 0.2838 0.2855 0.2915
SAGMAN 0.2808 0.2802 0.2808 0.2820 0.2814 0.2838
GOOD-AT+SAGMAN 0.2684 0.2695 0.2684 0.2690 0.2690 0.2701
@@ 4.2 Metrics for GNN Stability Evaluation
Figure 3: The horizontal axes, denoted by $X$, represent the perturbation applied. ‘Random EdgesPT’ refers to the DICE attack. ‘Random FeaturePT’ indicates the application of Gaussian noise perturbation, expressed as $X\eta$ perturbation, where $\eta$ represents Gaussian noise. SAGMAN Stable/Unstable denotes the samples classified as stable or unstable by SAGMAN, respectively.
We empirically demonstrate SAGMAN’s ability to distinguish between stable and unstable samples under DICE attack (Waniek et al., 2018) and Gaussian perturbation, as shown in Figure 3. Additional results on various datasets—including large-scale datasets—and different GNN architectures under Nettack attacks (Zügner et al., 2018) are provided in Appendix A.10, further illustrating the effectiveness of our metric.
@@ 4.3 Stability of GNN-based Recommendation Systems
We evaluate the stability of GNN-based recommendation systems using the PinSage framework (Ying et al., 2018) on the MovieLens 1M dataset (Harper & Konstan, 2015). To construct the input graph, we first homogenize the various node and edge types into a unified format. We then calculate the weighted spectral embedding matrix $U_{k}$ (as defined in Definition 3.1) and extract user-type samples to build a low-dimensional input (user) graph-based manifold. For the output graph-based manifold, we utilize PinSage’s final output, which includes the top-10 recommended items for each user. We construct this output (user) graph-based manifold based on Jaccard similarity measures of these recommendations. Table 6 demonstrates the effectiveness of SAGMAN in distinguishing between stable and unstable users.
Table 6: Comparison of the mean Jaccard similarity (MJS) between SAGMAN-selected stable/unstable users at a perturbation level ($l$) ranging from $1$ to $5$ where each selected user is connected to $l$ randomly chosen new items. The MJS is computed over $20$ iterations for each perturbation.
Perturbation Level 1 2 3 4 5
Stable Users 0.8513 0.8837 0.8295 0.8480 0.8473
Unstable Users 0.7885 0.7955 0.8167 0.8278 0.7794
@@ 4.4 SAGMAN-guided Adversarial Targeted Attack
We adapt our GNN training methodology from previous work (Jin et al., 2021). Using GCN as our base model for the Citeseer, Cora, Cora-ML, and Pubmed datasets, we employ Nettack and FGA (Chen et al., 2018) as benchmark attack methods. For target node selection, we compare Nettack’s recommendation (Zügner et al., 2018), SAGMAN-guided strategies, and a heuristic confidence ranking (Chang et al., 2017).
Table 3 presents the error rates after applying Nettack and FGA attacks. The results demonstrate that SAGMAN-guided attacks outperform both Nettack’s recommendation and the confidence ranking, leading to more effective adversarial attacks.